# Overcoming Knowledge Barriers: Online Imitation Learning from Observation with Pretrained World Models

### v1

@@ 3.2 Resolving the DKB
Based on the discussion from the previous sections, the straightforward way of solving the DKB is also to increase the number of demonstrations available to the agent. However, expert demonstrations are generally considered to be difficult and expensive to collect. Thus, increasing the size of the demonstration dataset is not always feasible in real-world applications.
@@
In order to propose a more practical solution, we need to look deeper into what is the real cause of the DKB. The policy learning part of the AIME algorithm is essentially behaviour cloning, and it is only conducted on the demonstration dataset. So for the states covered in the demonstration dataset, the policy is given clear guidance about what to do, while for other states, the behaviour is undefined. AIME is solely relying on the generalisation abilities of the learned latent state and the trained policy network to extrapolate the correct behaviour. In particular for small demonstration datasets, this can be unreliable or even impossible. Therefore, if we can enlarge the space of the covered states, we should reduce the DKB [51].
Generative Adversarial Imitation Learning (GAIL) style algorithms [27, 45, 58, 37] are perfect examples of this solution. Instead of directly learning from the demonstration, they adversarially train a discriminator to assess how how closely each state resembles those in the demonstration dataset. When learning the policy, they treat the discriminator’s output as a reward and encourage visiting states that are more likely belonging to the demonstrations. This modification provides guidance for newly visited states in the replay buffer, allowing the space of supported states to grow over the course of training.
However, this benefit is not clearly separable in GAIL since it is always entangled with the adversarial training of the discriminator. A recent work MAHALO [35] shows further evidence of the importance of the size of the covered space. The authors studied a similar ILfO setup with embodiment and demonstration datasets. They compared two variants: for one they train an inverse dynamics model (IDM) from the embodiment dataset and use it to label the demonstration dataset, while for another they train a reward model from the demonstration dataset by labelling all time steps with a reward of $1$, and then use it to label the embodiment dataset. Finally, they run the same offline RL algorithm on both labelled datasets. The results show the second variant attains a much better performance even though the labelling from the reward model is not as meaningful as the actions from the IDM.
Based on these insights, we propose to introduce a surrogate reward providing guidance signal for the agent on the replay buffer dataset. Due to the instability of adversarial training [17, 2, 27] and our focus on the pretraining paradigm, we opt to adopt the VIPER algorithm [14]. Instead of training a discriminator, VIPER trains a video prediction model on the demonstration datasets and treats the likelihood of the video prediction model as the reward for policy learning, i.e.
$r^{\mathrm{VIPER}}_{t}=\log p_{\nu}(o_{t}|o_{<t}).$ (5)
@@
Using this reward, we train the policy with a dreamer-style actor-critic algorithm [19] based on imagination in the latent space of the world model [19]. In order to do this, we first need to modify the reconstruction term in Equation 1 by adding an extra term for decoding the VIPER reward, i.e. $p_{\theta}(r^{\mathrm{VIPER}}_{t}|s_{t})$. Then, we further train a value estimator $V_{\xi}(s_{t})$ using TD($\lambda$)-return estimates, i.e.
@@
$\text{with}\;V^{(n)}_{\xi}(s_{t})$ $=\sum_{t^{\prime}=t+1}^{t+n}{\gamma^{t^{\prime}-t-1}r^{\mathrm{VIPER}}_{t^{\prime}}}+\gamma^{n}V_{\xi}(s_{t+n}).$
Using this estimate, we optimise our value function by minimising the MSE
@@
Using this value estimate, we extend the policy objective of Equation 3 to
@@
where $\beta$ is a hyper-parameter for balancing the two terms. We set $\beta=0.1$ by default based on the difference of default learning rate in AIME and Dreamer.
@@ 4 Experiments
We aim to answer the following questions: a) How does the proposed AIME-v2 compare with state-of-the-art methods on common benchmarks? b) How well does the proposed modification resolve the EKB and the DKB? c) How do different choices of hyper-parameters influence the results? In order to answer these questions, we design our experiments on DMC and Meta-World benchmarks.
4.1 Datasets
@@
For the DMC benchmark, we choose 9 tasks across 6 embodiments following Liu et al. [37] and use their published dataset as the demonstration datasets. Each dataset contains only 10 trajectories to reflect the scarcity of expert demonstrations. For the embodiment dataset, in order not to leak the task information from the pretraining phase, we follow Rajeswar et al. [48] and run a Plan2Explore [52] agent for each embodiment with 2M environments steps and use its replay buffer as the embodiment dataset. Different to them taking the model directly from the Plan2Explore agent as the pretrained model, we follow Zhang et al. [65] to retrain the model for $200\mathrm{k}$ gradient steps to get a better model.
@@
For Meta-World benchmark, we use the data and model from Hansen et al. [23]. The embodiment dataset was created from the replay buffer datasets. The open-sourced replay buffer datasets contain $40\mathrm{k}$ trajectories for each of the $50$ tasks with only state information. In order to fit to our image observation setup, we render the images by resetting the environment to the initial state of each trajectory and then executing the action sequence. The details of resetting to the initial state can be found in Appendix D.
@@
Following the idea of not leaking too much about the task information, inspired by the common practice in offline RL benchmarks [16], we use the first $200$ trajectories from each replay buffer and form a dataset with $10\mathrm{k}$ trajectories in total. We call this dataset MW-mt50. To further study the out-of-distribution transfer ability of the pretrained model, we follow the difficulty classification of the tasks from [53] and only use the $39$ easy and medium tasks to generate the datasets and the $11$ tasks hard and very hard tasks as hold-out tasks. We uniformly sample $250$ trajectories from te first $10\mathrm{k}$ trajectories from each of the $39$ tasks and form a dataset with $9750$ trajectories in total. We refer to this dataset as MW-mt39. Hence, MW-mt39 contains some expert behaviour solving the tasks, while MW-mt50 consists of mostly exploratory behaviour.
As the demonstration datasets, we use the single-task policies open-sourced by TD-MPC2 and collect $50$ trajectories for each tasks. We ensure that every trajectory in the demonstration dataset is successful. Since there are $500$ steps in a DMC trajectory and only $100$ steps in a Meta-World trajectory, the resulting datasets are roughly the same size.
Figure 2: Benchmark results and 9 DMC tasks. Return are calculated by running the policy 10 times with the environment and taking the average return. The results are averaged across 3 seeds with the shade region representing 95% CI.
@@ 4.2 Implementation
For the world model, we use the RSSM architecture [20] with the hyper-parameters in Hafner et al. [19] for DMC tasks. In addition, we use the KL Balancing trick from Hafner et al. [21] to make the training more stable. For Meta-World, since the visual scene is more complex, we use the M size model from Hafner et al. [22], but still with the continuous latent variable to be aligned with other models used in this paper. The policy network is implemented with a two-layer MLP, with $128$ neurons for each hidden layer. All the models are trained with Adam optimiser [31]. More details about the hyper-parameters can be found in Appendix C.
For the VIPER model, in the original paper, the authors first pretrain a VQ-GAN [15] and then train a GPT-style auto-regressive model in the quantised space for prediction. For simplicity of the implementation, in this paper, we consider training an unconditioned latent world model as in Seo et al. [54] to model the VIPER reward. We use the same RSSM architecture of the model learning for DMC, only removing the condition of the actions, and we train the VIPER model for each task separately. Especially during training, we find training such a powerful model from scratch on a small dataset can easily result in over-fitting. Thus, we empirically choose to train the model only for $500$ gradient steps for DMC models and $1000$ gradient steps for Meta-World models. We show evidence of overfitting in Appendix E. Due to the large scale of the ELBO, we also apply symlog [22] when computing the VIPER reward. Another difference with the original VIPER paper is that we do not use intrinsic motivation as the exploration bonus as the authors suggested, since the AIME loss for policy learning already provides task-related guidance for exploration. We only apply an entropy regulariser to the policy as is common practice. We further show the synergy between AIME and VIPER in Appendix F.
4.3 Benchmark Results
Figure 3: Benchmark results and $6$ Meta-World tasks. Trajectories are only counted as success when it success at the last time steps and the success rates are calculated with $10$ policy rollouts. The results are averaged across $3$ seeds with the shade region representing $95$% CI.
The benchmark results of DMC are shown in Figure 2. AIME-v2 significantly outperforms the PatchAIL baseline in $7$ out of $9$ tasks in terms of sample efficiency. Benefiting from the pretrained world model, AIME-v2 typically can reach expert performance within $200$k environment steps. Compared with BCO($\alpha$), updating the model is regularised and is benefitting more from the online interaction to resolve the EKB. Compared with the original offline AIME, AIME-v2 reliably improves performance, especially in hard tasks such as walker-run and hopper where offline AIME did not manage to make any progress.
However, there are still two tasks for which AIME-v2 does not show much progress, namely cartpole-swingup and quadruped-run. For cartpole-swingup, we observe that the policy learns to move the cart out of the scene so that the static image yields a high likelihood from the video prediction model. A similar phenomenon was also discussed in the original VIPER paper [14]. For quadruped-run, we conjecture that it is due to visual difficulties of a reconstruction-based model. When the quadruped is initialised on the ground, due to the symmetric structure of the robot, it is impossible to figure out which action corresponds to which leg, and it easily leads the action inference process to diverge. We additionally show AIME-v2 can work on these tasks with the help of the true reward in Appendix F.
The benchmark results of Meta-World are shown in Figure 3. We choose 4 tasks, namely disassemble, assembly, hand-insert and push from hard and very hard difficulties, and 2 tasks, namely sweep and hammer from the medium difficulty. While PatchAIL does not work on these tasks at all, AIME and AIME-v2 can make progress on them. AIME-v2 with either pretrained model outperforms AIME in all the tasks. On hard and very hard tasks, AIME with the mt50 model is better than AIME with the mt39 model. This is because they contain an unseen novel object by the mt39 model creating a large EKB. But in the online setting of AIME-v2, the two models are mostly on par. Moreover, using the mt50 models is better than using the mt39 models on average, which may imply covering diverse bahaviour is more valuable than knowing the expert directly.
4.4 Ablation Results
(a) Demonstration efficiency.
(b) Ablation of $\alpha$.
(c) Ablation of $\beta$.
@@ 4.4 Ablation Results
Figure 4: (a) Performance of AIME-v2, MBBC, AIME with respect to different number of demonstrations. For AIME-v2, we do not show the result for more than $100$ demonstrations since it is already saturated to the expert. (b) Ablation for the choice of the regulariser ratio $\alpha$. The left figure shows the mean return over $10$ trajectories while the right figure shows the MSE between the inferred actions and the true actions. (c) Ablation for the choice of the weight of the value gradient loss $\beta$. All results are averaged across $3$ seeds with the shaded region representing a $95$% CI.
We conduct our ablation studies on walker-run task from DMC.
How well does the proposed methods resolve knowledge barriers? In order to show how well AIME-v2 resolves the two knowledge barriers, we conduct a similar experiment as Zhang et al. [65] by providing the agent with different numbers of demonstrations. The result is shown in Figure 4(a). As we discussed before, MBBC as an oracle method that circumvents the EKB is a strict upper bound for AIME. And AIME-v2 which addresses both the EKB and DKB achieves much better results and is an upper bound for MBBC. From AIME-v2 can achieve near-expert performance with as few as $5$ demonstrations for this challenging task.
Influence of the data regulariser ratio $\alpha$. We set the regulariser ratio $\alpha$ from $[0.0,0.25,0.5,0.75]$ and plot the results in Figure 4(b). As we can see from the result, as long as we enable the regulariser, i.e. set $\alpha>0$, we get reliable improvements over the course of training. But if we disable the regulariser by setting $\alpha=0$, the learning only starts once there is enough new data in the replay buffer. As we discussed in Section 3.1, without the regulariser, in the early stage of the training, the model can easily overfit to the replay buffer, and it explains the early flattening phase of the training. As the training progresses, more and more data is available from the replay buffer, and it can establish the regulariser on its own, which explains the dramatic growing phase of the curves.
We also plot the MSE between the inferred actions and the true actions during the training process. From that we can see that a higher regulariser ratio offers more stable inference of the actions in the early phase of training.
@@
Influence of the value gradient loss weight $\beta$. We set the weight $\beta$ from $[0.0,0.01,0.05,0.1,0.5,1.0]$ and plot the results in Figure 4(c). As the result shows, having a small $\beta$ slows learning progress toward convergence. On the other hand, setting $\beta$ to a much larger value will improve the sample-efficiency without causing instability. For the sample efficiency, since we only have $10$ demonstrations, DKB dominates over EKB as shown in Figure 4(a). Thus, having a larger $\beta$ will make the learning much faster. In terms of the stability, as we discussed in 3.2, AIME loss and the value gradient loss operate on different regions of the environment states. This could make their influence on the policy independent of each other.

### v2

@@ 3.2 Overcoming the DKB
Based on the discussion from the previous sections, the straightforward way of overcoming the DKB is also to increase the number of demonstrations available to the agent. However, expert demonstrations are difficult and expensive to collect. Increasing the size of the demonstration dataset is not always feasible in real-world applications. Alternatively, we can modify the policy learning objective $J_{\mathrm{policy}}$ in Equation 3 to reduce the generalisation gap. In order to propose a practical solution, we first need to look deeper into what is the real cause of the DKB.
@@
The policy-learning part of the AIME algorithm is essentially behaviour cloning, and it is only conducted on the demonstration dataset. So for the states covered in the demonstration dataset, the policy is given clear guidance about what to do, while for other states, the behaviour is undefined. AIME solely relies on the generalisation abilities of the learned latent state and the trained policy network to extrapolate the correct behaviour. In particular for small demonstration datasets, this can be unreliable or even impossible. Therefore, if we were able to enlarge the space of the covered states, we should reduce the DKB (Ross et al., 2011).
@@
Based on these insights, we propose to introduce a surrogate reward providing a guiding signal for the agent on the replay buffer dataset, i.e. $r^{\mathrm{sur}}_{0:T}=R_{\nu}(o_{0:T})$. Using this reward, we train the policy with a dreamer-style actor-critic algorithm based on imagination in the latent space of the world model (Hafner et al., 2019a). In order to do this, we first need to modify the reconstruction term in Equation 1 by adding an extra term for decoding the surrogate reward, i.e. $\log p_{\theta}(r^{\mathrm{sur}}_{t}|s_{t})$. Then, we further train a value estimator $V_{\xi}(s_{t})$ using TD($\lambda$)-return estimates, i.e.
@@
$\text{with}\;V^{(n)}_{\xi}(s_{t})$ $=\sum_{t^{\prime}=t+1}^{t+n}{\gamma^{t^{\prime}-t-1}r^{\mathrm{sur}}_{t^{\prime}}}+\gamma^{n}V_{\xi}(s_{t+n}).$
Using this estimate, we optimise our value function by minimising the MSE, i.e.
@@
Using this value estimate, we extend the policy-learning objective $J_{\mathrm{policy}}$ of Equation 3 to
@@
where $\beta$ is a hyper-parameter for balancing the two terms. We set $\beta=1.0$ by default in this paper.
There could be many choices to derive this surrogate reward model. In this paper, we consider three different types of surrogate reward, namely AIL, OT and VIPER. AIL (Ho & Ermon, 2016; Torabi et al., 2019) uses adversarial training to learn a discriminator to tell whether an observations is generated by the expert, and uses the score from the discriminator as the reward. OT (Papagiannis & Li, 2023; Haldar et al., 2022) uses optimal transport theory to measure the distance between a given trajectory and a group of expert trajectories, and uses the negative distance as the reward. VIPER (Escontrela et al., 2023) learns a video prediction model from the demonstration datasets and uses the likelihood from the trained model as the reward. The detailed explanation of the three variants are in Appendix A. We use AIL as the default variant for AIME-NoB.
@@ 4 Experiments
In the experiments, we aim to answer the following questions: Q1: How does the proposed AIME-NoB compare with state-of-the-art methods on common benchmarks? Q2: How well do the proposed modifications overcome the EKB and the DKB? Q3: How do different design choices, components and hyper-parameters influence the results? In order to answer these questions, we design our experiments on DMC and MetaWorld benchmarks. The main results and the setup are described in the following sections, while some additional interesting results are presents in Appendix I.
4.1 Datasets and Tasks
@@
For the DMC benchmark, we choose nine tasks across six embodiments following Liu et al. (2022a) and use the same published dataset (Haldar et al., 2022) as the demonstration datasets. Each dataset contains only 10 trajectories to reflect the scarcity of expert demonstrations. For the embodiment dataset, in order not to leak the task information from the pretraining phase, we follow Rajeswar et al. (2023) and run a Plan2Explore (Sekar et al., 2020) agent for each embodiment with 2M environments steps and use its replay buffer as the embodiment dataset. Different to them taking the model directly from the Plan2Explore agent as the pretrained model, we follow Zhang et al. (2023) to retrain the model for 200k gradient steps to get a better model. When evaluating the performance of the learned policy on each task, we rollout the policy 10 times with the environment, and report the mean return.
@@
For vision-based MetaWorld benchmark, we use the data from Hansen et al. (2023a). The embodiment dataset was created from the replay buffer datasets. The open-sourced replay buffer datasets contain 40k trajectories for each of the 50 tasks with only state information. In order to fit to our image observation setup, we render the images by resetting the environment to the initial state of each trajectory and then executing the action sequence. The details can be found in Appendix E.
@@
With respect to the embodiment dataset, following the idea of not leaking too much about the task information, inspired by the common practice in offline RL benchmarks (Fu et al., 2021), we use the first 200 trajectories from each replay buffer and form a dataset with 10k trajectories in total. We call this dataset MW-mt50 and we use it for the benchmark on MetaWorld to compare AIME-NoB with other algorithms. To further study the out-of-distribution transfer ability of the pretrained model, we follow the difficulty classification of the tasks from (Seo et al., 2022a) and only use the 39 easy and medium difficulty tasks to generate the datasets and use the 11 hard and very hard tasks as hold-out tasks. We uniformly sample 250 trajectories from the first 10k trajectories from each of the 39 tasks and form a dataset with 9750 trajectories in total. We refer to this dataset as MW-mt39.
For evaluating the algorithms, we choose four hard or very hard tasks, namely disassemble, assembly, hand-insert and push; and two medium difficult tasks, namely sweep and hammer. As for the demonstration datasets, we use the single-task policies open-sourced by TD-MPC2 and collect 50 trajectories for each task. We ensure that every trajectory in the demonstration dataset is successful. For evaluation on each MetaWorld task, due to the noisy nature of the task, we rollout the policy 100 times with the environments, and report success only if the very last time step of an episode is marked as successful by the environment (following Hansen et al. (2023a)).
@@ 4.2 Implementation
For the world model, we use the RSSM architecture (Hafner et al., 2019b) with the hyper-parameters in Hafner et al. (2019a) for DMC tasks. In addition, we use the KL Balancing trick from Hafner et al. (2020) to make the training more stable. For MetaWorld, since the visual scene is more complex, we use the M size model from Hafner et al. (2023), but still with the continuous latent variable to be aligned with other models used in this paper. The policy network is implemented with a two-layer MLP, with 128 neurons for each hidden layer. All the models are trained with Adam optimiser (Kingma & Ba, 2017). More details on hyper-parameters are in Appendix C.
Figure 2: Comparing AIME-NoB with other algorithms. The figures show aggregated IQM scores on 9 DMC tasks and 6 MetaWorld tasks. All the algorithms are evaluated with 5 seeds on each task and the shaded region representing 95% CI.
4.3 Results
Q1: How does AIME-NoB compare with state-of-the-art methods on common benchmarks? We compare the AIL variant of AIME-NoB with the representative state-of-the-art algorithms:
  • AIME (Zhang et al., 2023) represents the base algorithm that we improve upon. The algorithm uses pretrained world model offline, and suffers from both EKB and DKB.
  • BCO($\alpha$) (Torabi et al., 2018) is an online extension of the popular BCO(0) algorithm, which represents another line of methods that can use pretrained models. The online interactions in BCO($\alpha$) can potentially overcome EKB from BCO(0).
  • PatchAIL (Liu et al., 2022a) represents the Generative Adversarial Imitation Learning (GAIL) styles of algorithms (Ho & Ermon, 2016; Torabi et al., 2019).
  • OT (Haldar et al., 2022) represents the trajectory matching based algorithms.
The results of the benchmark are shown in Figure 2. We follow Agarwal et al. (2022) to report the IQM score and 95% CI when aggregating over all the tasks in the same suite. For each DMC task, the score is normalised using the average return of the expert, while for each MetaWorld task we directly report the success rate. From the results, we can clearly see AIME-NoB achieves better sample-efficiency and final performance on both of the benchmark suites comparing with all other algorithms. Benefiting from the pretrained world model, AIME-NoB typically can reach near expert performance within 100k environment steps on DMC tasks. This even matches the performance in Rajeswar et al. (2023) where the true rewards are available. For the complete results of each individual task please refer to Appendix H.
@@ 4.3 Results
Figure 3: Performance of AIME-NoB, AIME-NoEKB, MBBC, AIME w.r.t. different number of demonstrations on the walker-run task. For AIME-NoB, we do not show the result for more than 20 demonstrations since it is already saturated to the expert. All results are averaged across 5 seeds with the shaded region representing a 95% CI.
Q2: How well do the proposed methods overcome knowledge barriers? In order to show how well AIME-NoB overcomes the two knowledge barriers, we conduct the same experiment as in Figure 1 by providing the agent with different numbers of demonstrations on walker-run. We also run an additional variant coined AIME-NoEKB where we only apply the solution for the EKB. The result is shown in Figure 3. As we discussed before, MBBC as an oracle method that circumvents the EKB is an upper bound for AIME. AIME-NoEKB matches and even slightly outperforms MBBC, which implies the proposed solution completely overcomes the EKB. The fact that it slightly outperforms MBBC is a bonus of the model choice – fine-tuning the latent variable world model improves the generalization of the latent space which mitigates the DKB. AIME-NoB, which further addresses the DKB, matches the expert performance even when given only 1 demonstration. This showcases that the DKB has also been completely overcome using the surrogate reward. The difference between different number of demonstrations is mainly on the sample-efficiency side, where the more demonstrations we have the less online interactions we need to attain the expert performance.
(a) Different variants on DMC.
(b) Different variants on MetaWorld.
(c) Different datasets on MetaWorld.
Figure 4: Ablations of different variants of AIME-NoB and choices of the embodiment datasets. All the algorithms are evaluated with 5 seeds with the shaded region representing 95% CI.
Q3.1: Which variant of AIME-NoB performs the best? We run all the three AIME-NoB variants with different surrogate rewards, i.e. AIL, OT and VIPER, on the two benchmark suites and aggregate the results. From Figure 4(a) and 4(b), we can see the AIL variant of AIME-NoB generally performs the best on both suites. On the DMC suite, all the three variants managed to converge to the same final performance, but the AIL variant is slightly more sample efficient. On the MetaWorld suite, the distinction is larger between different variants, which highlights the supremacy of the AIL variant. We hypothesise that AIL is the only variant that directly adapts the surrogate rewards online by training the discriminator. For the other two variants: VIPER has zero adaptivity since the VIPER model is fixed during the online training; OT has an indirect adaptivity since it depends on the image encoder from the world model which get finetuned during online learning. This adaptivity enables AIL to capture more pertinent signals during the imitation process. Based on these results, we choose the AIL variant as the default implementation for AIME-NoB.
Q3.2: Which dataset can pretrain the better world model for AIME-NoB? The quality of the pretrained models naturally depends on the quality of the pretraining datasets. In order to understand what characteristics of the datasets influence the performance, we train all the AIME-NoB variants with two different models pretrained separately on the MW-mt39 datasets and MW-mt50 datasets. The aggregated result with each dataset is shown in Figure 4(c). The result demonstrates that although the size of the datasets is roughly the same, the model pretrained on MW-mt50 offers better results. This may imply covering diverse behaviours and objects is more valuable than knowing the expert, for example the mt50 dataset contains objects to be assembled while the mt39 does not.
Figure 5: Ablations of replay ratio $\alpha$ on walker-run task. AIME-NoB is running with 10 demonstrations, while AIME-NoEKB is with 100 demonstrations. Action MSE is only shown for the first $10^{5}$ env steps. All results are averaged across 5 seeds with the shaded region representing a 95% CI.
Q3.3: How does different data regulariser ratio $\alpha$ influence the performance? We ablate the regulariser ratio $\alpha$ from [0.0, 0.25, 0.5, 0.75]. Further, we compare to a simple append version where the online dataset is appended to the embodiment dataset and treated as a singular dataset for sampling. The append version can be also understood in this experiment as having an inverse proportional schedule of $\alpha$ from 1.0 to 0.66 during the course of training. To isolate the effect on the EKB, we train both AIME-NoB and AIME-NoEKB. We show both the training curves of the return and action MSE, which is the MSE between the inferred actions and the true actions, in Figure 5. For AIME-NoEKB, as long as we enable the regulariser, i.e. set $\alpha>0$, we get reliable improvements of returns over the course of training. But if we disable the regulariser by setting $\alpha=0$, the return exhibits high variance. The cause is clear by looking at the action MSE. For $\alpha=0$ the action MSE diverges in the beginning and cannot recover. For AIME-NoB, the story is more complicated. While the action MSE still diverges in the beginning when $\alpha\leq 0.5$, the surrogate reward can guide the policy back on track and even achieves lower action MSE after recovery. In this case, a small $\alpha$ helps the algorithm make more use of the online dataset, resulting in higher sample efficiency. Although we keep the choice of $\alpha$ with a fixed value for AIME-NoB in the light of simplicity, there should be a smart scheduling for each task to strike the balance between stability and sample-efficiency. However, we think finding this optimal strategy is beyond the scope of this paper.
(a) Effect of the pretrained world model. -m means removing the pretrained world model while -d means removing the pretrained dataset.
(b) Ablation of removing AIME loss from AIME-NoB. -p means removing from pretraining while -f means removing from finetuning.
(c) Ablation for the choice of the weight of the value gradient loss $\beta$.
Figure 6: Ablations of pretrained models, AIME loss and the weight of the value gradient loss $\beta$. All the algorithms are evaluated with 5 seeds with the shaded region representing 95% CI.
Q3.4: How much benefit do we get from the pretrained world model and the dataset for pretraining? One of the advantages of AIME-NoB over popular ILfO algorithms is that it can make use of pretrained models and pre-collected datasets. Thus, we want to investigate how much AIME-NoB benefits from having a pretrained model and a pre-collected dataset. We rerun AIME-NoB on walker-run without the embodiment dataset and without the pretrained world model. As we can see the result from Figure 6(a), without the pretrained the world model, the sample-efficiency is largely affected. Between the two components, the model is more important than the dataset.
@@
Q3.5: How do different values of the gradient loss weight $\beta$ influence the performance? We set the weight $\beta$ from [0.0, 0.1, 1.0, 10.0, 100.0] and plot the results in Figure 6(c). As the result shows, without the surrogate reward, i.e. $\beta=0$, the agent cannot reach expert performance due to the DKB. Having a small $\beta$ slows learning progress toward convergence. On the other hand, setting $\beta$ to a much larger value will improve the sample-efficiency without causing instability. For the sample efficiency, since we only have 10 demonstrations, the DKB dominates over the EKB as shown in Figure 3. Thus, having a larger $\beta$ will speed up learning. In terms of stability, as we discussed in 3.2, AIME loss and the value gradient loss operate on different regions of the environment states. This could make their influence on the policy independent of each other.
Q3.6: Is surrogate reward all you need? Given the results above that AIME-NoB can work well even when we lower the effect of the AIME loss by making either worse action inference, i.e. set $\alpha$ to low value, or strengthen the value gradient loss, i.e. set $\beta$ to high value, a natural question to ask is whether the AIME loss is still needed in AIME-NoB or whether surrogate reward only is enough to solve the imitation learning task. The AIME loss is used in two places in the AIME-NoB algorithm – both for pretraining the policy offline and finetuning the policy online. We compare AIME-NoB with variants that removes AIME from these parts. Form the results shown in Figure 6(b), removing AIME loss from either pretraining or finetuning will lower the sample efficiency. Removing from both phases causes instability and convergence issues during training. Thus, AIME loss is still crucial and cannot simply be replaced by surrogate rewards.